OpenStudy (anonymous):

What is the simplified version of (6x^2y^3)(4xy^3)?

OpenStudy (anonymous):

Is it 24x^3y^6

OpenStudy (anonymous):

Or 24x^3y^9

OpenStudy (campbell_st):

well you will use the index rule for multiplying \[x^a \times x^b = x^{a + b}\] so its \[6 \times 4 \times x^2 \times x \times y^3 \times y^3\]

OpenStudy (campbell_st):

your first answer is correct